The yield y(t) (in bushes)per acre of a corn crop satisfies the equation dy/dt+y=100+e-t

if y(0)=0 find y at any time t



1
Expert's answer
2021-07-07T11:21:47-0400
"y'+y=100+e^{-t}"

Integrating factor: "\mu=e^t"


"(e^ty)'=100e^t+1"

"e^ty=\int(100e^t+1)dt"

"e^ty=100e^t+t+c"

"y(t)=100+te^{-t}+ce^{-t}"

"y(0)=0"


"0=100+0+ce^{-0}=>c=-100"


"y(t)=100+te^{-t}-100e^{-t}"